OpenStudy (anonymous):

An evacuated glass vessel weighs 50.0 g when empty,148.0 g when filled with a liquid of density 0.98 g/ml and 50.5 g when filled with an ideal gas at 760 mm Hg at 300 K. Determine the molecular mass of the gas.

OpenStudy (anonymous):

volume of glass vessel=weight of liquid divided by density=100 ml. using ideal gas equation and changing all units to litre and atmosphere. pv=nrt, so 1*0.1=n*0.0821*300, therefore n=0.00406. and n= weight/molar mass, so molar mass=0.5/0.00406=123.153 .